if a certain number of 2 digits is divided by the sum of its digits the quotient is 3 and remainder is 7 . If the digits are interchanged and the resulting number is divided by the sum of digits the quotient is 7 and the remainder is 3. Find the number.
Answer by dkppathak(439)

let unit place is y
and tens place is x of two digit number
number 10x+y
and reverse number will be
10y+x
as per given conditions
10x+y=3(x+y)+7
and
10y+x +7(x+y)+3
solving these equations
we can get as
7x-2y=7 i
-6x+3y=3 ii
by solving both equations elimination method
we will get
X=3
and Y=7
two digit number whose tens place is 3 and unit place is 7
two digit number is 37
answer 37